# Quantum Metrology via Floquet-Engineered Two-axis Twisting and Turn Dynamics

### v1

@@ Abstract
The core of quantum metrology lies in utilizing entanglement to enhance measurement precision beyond standard quantum limit. Here, we utilize the Floquet-engineered two-axis twisting (TAT) and turn dynamics to generate non-Gaussian states for quantum metrology. By employing both analytically semi-classical and quantum approaches, we find that the desired $N$-particle non-Gaussian state can be produced within a remarkably short time $t_{\mathrm{opt}}\propto\ln{N}/{N}$, and its quantum Fisher information $F^{\mathrm{opt}}_{\mathrm{Q}}\propto N^{2}$ approaches the Heisenberg limit. Moreover, using the Floquet-engineered anti-TAT-and-turn, we may implement an efficient interaction-based readout protocol to extract the signal encoded in this non-Gaussian state. This Floquet-engineered anti-TAT-and-turn approach offers a viable method to achieve effective time-reversal dynamics for improving measurement precision and resilience against detection noise, all without the need to invert the sign of the nonlinear interaction. This study paves the way for achieving entanglement-enhanced quantum metrology via rapid generation of cat-like states at high particle numbers through continuous Floquet engineering.
@@
Introduction.— Quantum metrology utilizes quantum entanglement to enhance measurement precision from the standard quantum limit (SQL) to the Heisenberg limit (HL) [1, 2, 3, 4, 5, 6, 7, 8], offering potential applications in atomic clocks [9, 10, 11, 12, 13, 14, 15, 16], magnetometers [17, 18, 19, 20, 21, 22, 23, 24], gyroscopes [25, 26, 27, 28], gravimeters [29, 30, 31] and other sensors. The ultimate measurement precision of an estimated phase $\phi$ is determined by the quantum Cramér-Rao bound (QCRB) $\Delta\phi\geq 1/\sqrt{F_{Q}}$, defined by the quantum Fisher information (QFI) ${F_{Q}}$  [32, 33, 34]. How to prepare an achievable entangled state with high QFI is a top priority. One-axis twisting (OAT) dynamics [35, 36] stands out as a crucial mechanism for generating entanglement. Using OAT dynamics, the Gaussian-shaped spin squeezed states have been successfully demonstrated in several proof-of-principle experiments [37, 38, 39, 40, 41, 42]. Although non-Gaussian entangled states [43, 44, 45, 46] (especially the spin cat states [47, 48, 49] or Greenberger-Horne-Zeilinger (GHZ) state [50, 51]) offer higher QFI [52, 53], they require significantly longer evolution times for preparation, presenting a substantial practical challenge.
@@
It has been demonstrated that OAT-and-turn [54, 55] and two-axis twisting (TAT) [56, 57, 58] can achieve faster entanglement generation. Although TAT exhibits a superior speedup, its interaction form does not naturally exist in known physical systems [59, 60]. In principle, an effective TAT interaction can be realized with OAT-and-turn, in which one can transform OAT into TAT by applying suitable transverse coupling [61, 62, 63, 64]. The coupling field can be a train of periodic pulses [61, 62], a periodically modulated continuous field [63], or even a sequence designed by machine learning [64]. Since the turn dynamics can accelerate OAT to achieve fast entanglement generation, a natural question arises: can it accelerate TAT to achieve faster entanglement generation?
@@
On the other hand, extracting the signal encoded in entangled states generally necessitates single-particle-resolution measurements of parity [51, 65], high-order observables [66, 67], or full probability distribution [44, 45]. However, the single-particle-resolution measurements are very susceptible to detection noise [68, 69]. To address this challenge, time-reversal interaction-based readout provides a powerful protocol for achieving high-precision Heisenberg-limited measurements [70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82]. This requires to reverse the dynamics of an interacting many-body quantum system, which is usually accomplished by inverting the sign of nonlinear interaction [73, 74, 75]. Can one realize time-reversal TAT-and-turn dynamics to achieve Heisenberg-limited measurement without changing the sign of nonlinear interaction?
@@
In this Letter, we employ Floquet-engineered TAT-and-turn dynamics to efficiently create non-Gaussian entangled states and carry out the necessary time-reversal readout. By applying a suitable transverse coupling field, an effective TAT-and-turn dynamics of an ensemble of Bose condensed two-level atoms can be realized via Floquet engineering. Using both analytically semi-classical and quantum treatments, we find that non-Gaussian entangled states with high QFI can be generated in much shorter time. Furthermore, the time-reversal readout can be realized by applying another Floquet-engineered TAT-and-turn dynamics without flipping the sign of nonlinear interaction. It allows to efficiently extract the signal encoded in the prepared non-Gaussian entangled states and thus provides a robust approach for enhancing measurement precision. Our work develops an efficient protocol for performing entanglement-enhanced quantum metrology via rapid generation and detection of cat-like states through continuous Floquet engineering.
@@
Figure 2: The maximum QFI of states generated by the Floquet-engineered TAT-and-turn dynamics. (a) The maximum QFI $F_{Q}^{\mathrm{max}}$ versus evolution time $\chi t$ under different $\delta$ with particle number $N=100$. The amplitude and initial phase of Rabi frequency are chosen as $\Omega_{0}=2\pi\times 100N\chi$ and $\alpha=0$, respectively. (b) The time evolution of maximum QFI for Floquet-engineered TAT-and-turn with optimal detuning $\delta_{\mathrm{opt}}$, and the results of OAT, OAT-and-turn, TAT, and TAT-and-turn are presented for comparison. (c) The probability distribution and Husimi distribution (inset) of the optimal state with the largest QFI in (b), where the probability distribution is projected onto the eigenstates of $\hat{J}_{y}$ with $P_{m_{y}}=|\langle m_{y}|\psi\rangle|^{2}$. (d) The critical detuning $\delta_{\mathrm{opt}}$ and (e) the critical evolution time $\chi t_{\mathrm{opt}}$ corresponding to (f) the optimal QFI $F_{Q}^{\mathrm{opt}}$ for different particle numbers $N$, where the semi-classical predictions are $\delta_{\mathrm{opt}}^{\mathrm{SC}}/(N\chi)\approx 0.3135$, $t_{\mathrm{opt}}^{\mathrm{SC}}={3(1.9+0.55\ln{N})}/{N}$ and $(F_{Q}^{\mathrm{opt}})^{\mathrm{SC}}/N\approx 0.97N$, respectively. The fitted line for the optimal $F_{Q}^{\mathrm{max}}$ is $F_{Q}^{\mathrm{opt}}/N\simeq 0.77N$.
@@
Then, we perform Fourier decomposition on this Hamiltonian. Using the Euler formula $e^{i\gamma}=\cos{\gamma}+i\sin{\gamma}$ and the Jacobi-Anger expansion $e^{iz\sin\theta}=\mathcal{{J}}_{0}(|z|)+\sum_{n=1}^{\infty}[\mathcal{{J}}_{n}(z)e^{in\theta}+(-1)^{n}\mathcal{{J}}_{n}(z)e^{-in\theta}]$ with $\mathcal{{J}}_{n}(z)$ being the $n$-th Bessel function of the first kind, the Hamiltonian (2) becomes
@@
where $\hat{H}_{0}^{I}=\frac{\chi}{2}[(1+L_{0})\hat{J}_{z}^{2}+(1-L_{0})\hat{J}_{\beta}^{2}]+K_{0}\delta\hat{J}_{z}$ and $\hat{H}_{n}^{I}$ represents $n$-order Hamiltonian with $L_{n}=\mathcal{{J}}_{n}\left(2\Omega_{0}/\omega\right)$ and $K_{n}=\mathcal{{J}}_{n}\left(\Omega_{0}/\omega\right)$ [88]. Finally, based on the Floquet-Magnus expansion [89, 90], up to the order of $\omega^{-1}$, the effective Hamiltonian for Eq. (11) can be written as $\hat{H}_{\mathrm{F}}^{\rm eff}=\hat{H}_{0}^{I}+\sum_{n=1}^{\infty}\frac{1}{n\omega}([\hat{H}_{n}^{I},\hat{H}_{-n}^{I}]-[\hat{H}_{n}^{I},\hat{H}_{0}^{I}]+[\hat{H}_{-n}^{I},\hat{H}_{0}^{I}])+\mathcal{O}(\omega^{-1})$. When $\omega$ is sufficiently large ($\omega\gg N\chi$), it reduces to $\hat{H}_{\mathrm{F}}^{\rm eff}\simeq\hat{H}_{0}^{I}$ [88]. For a conserved collective spin, $\hat{J}^{2}=\hat{J}_{\alpha}^{2}+\hat{J}_{\beta}^{2}+\hat{J}_{z}^{2}=\frac{N}{2}(\frac{N}{2}+1)$ is a constant, the effective time-independent Floquet Hamiltonian is equivalent to
@@
with the effective interaction strength $\chi_{\mathrm{eff}}=\chi/3$ and the effective detuning $\delta_{\mathrm{eff}}=K_{0}\delta\simeq 0.404\delta$. Since the TAT-and-turn Hamiltonian (5) is derived from the Hamiltonian (9) with the Floquet theorem, we refer the original time-evolution to the Floquet-engineered TAT-and-turn (labelled as FE-TATNT) dynamics. To evaluate the Floquet-engineered TAT-and-turn dynamics, we also calculate the dynamics of the ideal TAT-and-turn Hamiltonian (5) for comparison.
Varying system parameters, the Floquet-engineered TAT-and-turn dynamics can be reduced to typical TAT, OAT-and-turn, or OAT dynamics. When $\delta=0$, the effective TAT-and-turn Hamiltonian (5) can be reduced to an effective TAT Hamiltonian [56, 57, 58] $\hat{H}_{\rm TAT}=\chi_{\rm eff}(\hat{J}_{\beta}^{2}-\hat{J}_{\alpha}^{2})$. Further, setting $\Omega(t)=\Omega_{0}$ as a constant and $\delta=0$, the Hamiltonian (9) is reduced to an OAT-and-turn Hamiltonian [54, 55] $\hat{H}_{\mathrm{OATNT}}=\chi\hat{J}_{z}^{2}+\Omega_{0}\hat{J}_{\alpha}$. If $\delta=0$ and $\Omega(t)=0$, the Hamiltonian (9) becomes an OAT Hamiltonian $\hat{H}_{\mathrm{OAT}}=\chi\hat{J}_{z}^{2}$.
Characterization of full metrological potential for a given state.— The metrologically useful entangled states created by the above different schemes typically correspond to different optimal sensing directions. To evaluate the full metrological usefulness of the states generated by different schemes, we first quantify the best achievable measurement precision using the maximum QFI $F_{Q}^{\rm max}$ for all potential sensing orientations. For an arbitrary state, one can determine a $3\times 3$ quantum Fisher information matrix (QFIM) $\mathcal{F}$  [91] with matrix elements $\mathcal{F}_{\mu\nu}=2\langle\{\hat{J}_{\mu},\hat{J}_{\nu}\}\rangle_{\psi}-4\langle\hat{J}_{\mu}\rangle_{\psi}\langle\hat{J}_{\nu}\rangle_{\psi}$, where $\mu,\nu=x,y,z$. By solving the eigenequation $\mathcal{F}\vec{n}=\lambda\vec{n}$, one can obtain three eigenvalues $\lambda_{i}$ and their corresponding eigenvectors $\vec{n}_{i}$ (where $i=\{1,2,3\}$). The maximum eigenvalue $\lambda_{\rm max}$ and its eigenvector $\vec{n}_{\rm max}=(n_{x}^{\rm max},n_{y}^{\rm max},n_{z}^{\rm max})^{\mathrm{T}}$ correspond to the maximum QFI $F_{Q}^{\rm max}$ and the optimal sensing direction (whose optimal generator is $\hat{J}_{\vec{n}_{\rm max}}=n_{x}^{\rm max}\hat{J}_{x}+n_{y}^{\rm max}\hat{J}_{y}+n_{z}^{\rm max}\hat{J}_{z}$), respectively. Therefore, one can use $F_{Q}^{\rm max}$ to determine the full metrological potential for quantum sensing and rotate the state until the optimal generator $\hat{J}_{\vec{n}_{\rm max}}$ matches the Hamiltonian for a given sensing purpose [91].
@@
Fast generation of metrologically useful non-Gaussian entangled states with Floquet-engineered TAT-and-turn.— Using the above maximum QFI $F_{Q}^{\rm max}$, below we analyze the full metrological potential of the states generated by Floquet-engineered TAT-and-turn dynamics (9). Without loss of generality, we choose $\alpha=0$ hereafter. It is advantageous to begin with an initial spin coherent states on one of the poles along $z$-axis ($\lvert\uparrow\rangle^{\otimes N}$ or $\lvert\downarrow\rangle^{\otimes N}$), which correspond to two saddle points $(0,0,\pm 1)$ in the classical phase-space of TAT-and-turn when $\left\lvert\delta_{\rm eff}/\chi_{\rm eff}\right\lvert<N$, see Fig. 1 (c). To identify the optimal condition for Floquet-engineered TAT-and-turn, we analyze the dependence of $F_{Q}^{\rm max}$ on the detuning $\delta$ and the evolution time $\chi t$ in Fig. 2 (a). It clearly shows that the maximum QFI has a symmetrical distribution with respect to $\delta=0$. Given the particle number $N=100$, the maximum QFI reaches its optimal value $F_{Q}^{\mathrm{opt}}\simeq 0.7783N^{2}$ when the evolution time $\chi t_{\mathrm{opt}}\approx 0.132$ and the detuning $\delta_{\mathrm{opt}}\approx\pm 0.3135N\chi$. In Fig. 2 (b), we show the time evolution of $F_{Q}^{\rm max}$ for OAT, OAT-and-turn, TAT, TAT-and-turn ($\hat{H}_{\rm TATNT}$ with $\delta_{\mathrm{eff}}=K_{0}\delta_{\mathrm{opt}}$), and Floquet-engineered TAT-and-turn ($\hat{H}_{\rm FE}$ with $\delta=\delta_{\mathrm{opt}}$). The results of Floquet-engineered TAT-and-turn are well consistent with the ideal TAT-and-turn, thereby confirming the validity of the effective TAT-and-turn Hamiltonian (5). Compared with other schemes, the states generated via Floquet-engineered TAT-and-turn dynamics may achieve the largest $F_{Q}^{\rm max}$ when $0.10\lesssim\chi t\lesssim 0.18$. In terms of the best achievable QFI, the Floquet-engineered TAT-and-turn outperforms the OAT by approximately $54\%$, OAT-and-turn and TAT by around $20\%$.
Under the Floquet-engineered TAT-and-turn with $\chi t_{\mathrm{opt}}$ and $\delta_{\mathrm{opt}}$, the generated state is a non-Gaussian entangled state of two main parts distributed near the poles of the $y$-axis. Such a non-Gaussian state is similar to a spin cat state [47, 48, 49], see Fig. 2 (c). Moreover, the best achievable QFI of Floquet-engineered TAT-and-turn exhibits the Heisenberg scaling $F_{Q}^{\rm max}\propto N^{2}$, which is consistent with the QCRB of a spin cat state. Correspondingly, the evolution time for generating the state obeys $\chi t_{\rm opt}\propto\ln{N}/{N}$. This indicates that the required time can be very short for large particle number $N$, far less than that of the OAT dynamics, which is of essential importance and feasibility for overcoming decoherence.
@@
The Heisenberg scaling, the critical detuning and the critical evolution time can be analytically found by using semi-classical treatment [54, 92]. Since there is a special symmetry in the classical phase space of the TAT-and-turn Hamiltonian (5), the states generated by Floquet-engineered TAT-and-turn dynamics always have symmetric probability distributions. Just like the resembled spin cat state in the inset of Fig. 2 (c), whose variance is approximately stretched along a certain axis, one can estimate the global optimal parameter $\delta_{\rm opt}^{\rm SC}=\pm(\sqrt{2}-1)N\chi/(3K_{0})\approx\pm 0.3135N\chi$ by the upper boundary of the QFI from the Heisenberg limit [34]. Due to the center of the uncertainty patch is fixed, the TAT-and-turn dynamics merely alters the distribution by stretching and squeezing. Thus, the relevant timescale can be estimated by the time interval between the boundary points of the uncertainty patch along the separatrix [92], which can be given as [88]
@@
Based on the above analysis, the best achievable QFI exhibits the Heisenberg scaling, which matches the ultimate precision bound of spin cat states [88]. As shown in Fig. 2 (d-f), the above results are in accordance with the numerical results of Floquet-engineered TAT-and-turn. According to Eq. (6), the Floquet-engineered TAT-and-turn dynamics exhibits significantly shorter timescale for large particle numbers compared to the OAT and it is even faster than the GHZ-like state generation with effective three-body interaction [93]. Thus our protocol is advantageous in mitigating decoherence in practical quantum metrology experiments, showing remarkable metrological usefulness.
@@
Figure 3: (a) Measurement precision yielded from Floquet-engineered anti-TAT-and-turn. The results of time-reversal readout with OAT, OAT-and-turn, TAT and TAT-and-turn are presented for comparison. The QCRB of Floquet-engineered TAT-and-turn is also shown (green circle). (b) Robustness against detection noise. The metrological gain is defined as $G=20\log_{10}\left[{(\Delta\phi)_{\mathrm{SQL}}}/{\Delta\phi}\right]$ with $(\Delta\phi)_{\mathrm{SQL}}=1/{\sqrt{N}}$ denoting the measurement precision of SQL. Green dashed line is the result of optimal measurement $\hat{J}_{\vec{m}_{\rm max}}$ after Floquet-engineered anti-TAT-and-turn with $\chi t=0.12$. Green dashed-dot and yellow dot lines are the results with direct parity measurement for non-Gaussian state generated by Floquet-engineered TAT-and-turn at $\chi t=0.12$ and the GHZ state generated by OAT at $\chi t=\pi/2$, respectively. Here, the amplitude and initial phase of Rabi frequency are chosen as $\Omega_{0}=2\pi\times 50N^{2}\chi$ and $\alpha=0$ with particle number $N=100$ and the estimated phase $\phi=1/1000$.

### v2

@@ Abstract
One core of quantum metrology is the utilization of entanglement to enhance measurement precision beyond the standard quantum limit. Here, we utilize the Floquet-engineered two-axis twisting (TAT) and turn dynamics to generate GHZ-like states for quantum metrology. Using both analytical semi-classical and quantum approaches, we find that the desired $N$-particle GHZ-like state can be produced in a remarkably short time $t_{\mathrm{opt}}\propto\ln{N}/{N}$, and its quantum Fisher information $F^{\mathrm{opt}}_{\mathrm{Q}}\propto N^{2}$ approaches the Heisenberg limit. Owing to the rapid state preparation, it shows outstanding robustness against decoherence. Moreover, using the Floquet-engineered anti-TAT-and-turn, one may implement an efficient interaction-based readout protocol to extract the signal encoded in this GHZ-like state. This Floquet-engineered anti-TAT-and-turn approach offers a viable method to achieve effective time-reversal dynamics to improve measurement precision and resilience against detection noise, all without the need to invert the sign of the nonlinear interaction. This study paves a way for achieving entanglement-enhanced quantum metrology via rapid generation of GHZ-like states at high particle numbers through continuous Floquet engineering.
@@
Introduction.— Quantum metrology utilizes quantum entanglement to enhance measurement precision from the standard quantum limit (SQL) to the Heisenberg limit (HL) [1, 2, 3, 4, 5], offering promising applications in atomic clocks [6, 7, 8, 9], magnetometers [10, 11, 12, 13], gyroscopes [14, 15, 16, 17], gravimeters [18, 19, 20] and other sensors. The ultimate measurement precision of an estimated phase $\phi$ is determined by the quantum Cramér-Rao bound (QCRB) $\Delta\phi\geq 1/\sqrt{F_{Q}}$, defined by the quantum Fisher information (QFI) ${F_{Q}}$  [21, 22, 23]. The preparation of an achievable entangled state with high QFI is a top priority. Among all classes of entangled states, multi-particle maximally entangled state (i.e., Greenberger-Horne-Zeilinger (GHZ) state) [24, 25, 26, 27] and GHZ-like states that share similarities with the GHZ state [28, 29, 30, 31] (e.g., spin cat states [32, 97, 34]) are among the ultimate goals for both quantum metrology and quantum information [35]. Although the dynamics of the one-axis twisting (OAT) [36, 37] can generate these entangled states, it requires a significantly long evolution time for preparation[38], which poses a substantial challenge [39].
@@
It has been demonstrated that OAT-and-turn [102, 103] and two-axis twisting (TAT) [100, 43, 101] can achieve faster entanglement generation. Although TAT exhibits a superior speedup, its interaction form does not naturally exist in known physical systems [45, 46]. In principle, an effective TAT interaction can be realized with OAT-and-turn, in which one can transform OAT into TAT by applying suitable transverse couplings [47, 48, 49, 50]. The coupling field can be a train of periodic pulses [47, 48], a continuous periodic modulated field [49], or even a sequence designed by machine learning [50]. Besides, by applying Floquet-engineered pulses in OAT, an artificial three-body interactions (named as XYZ model) may be realized to rapidly generate GHZ-like states with high QFI [105]. A natural question arises: can we use turn dynamics to accelerate TAT and achieve faster GHZ-like state generation by applying Floquet-engineering?
@@
On the other hand, extracting the signal encoded in GHZ-like states generally requires single-particle-resolution measurements of parity [25, 52], high-order observables [53, 54], or full probability distribution [29, 30]. However, single-particle-resolution measurements are highly susceptible to detection noises [55, 56]. To address this challenge, time-reversal interaction-based readout provides a powerful protocol to achieve high-precision Heisenberg-limited measurements [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 104]. This requires reversing the dynamics of an interacting many-body quantum system, which is usually achieved by inverting the sign of nonlinear interaction [60, 61, 62]. Can one realize time-reversal dynamics to achieve Heisenberg-limited measurement without changing the sign of nonlinear interaction?
@@
In this Letter, we employ Floquet-engineered TAT-and-turn dynamics to efficiently create GHZ-like states and carry out the time-reversal readout without the need to invert the sign of the nonlinear interaction. By applying a suitable transverse coupling field, an effective TAT-and-turn dynamics in an ensemble of Bose condensed two-level atoms can be realized via Floquet engineering. Using both analytical semi-classical and quantum treatments, we find that GHZ-like states with high QFI can be generated in a much shorter time which is advantageous in mitigating the effect of decoherence. Furthermore, the time-reversal readout can be realized by applying another Floquet-engineered TAT-and-turn dynamics without flipping the sign of nonlinear interaction, which is experimentally feasible. It allows us to efficiently extract the signal encoded in the prepared GHZ-like state, and thus provides a robust approach for enhancing measurement precision. Our work develops an efficient protocol for performing entanglement-enhanced quantum metrology via rapid generation and detection of GHZ-like states through continuous Floquet engineering.
@@
Figure 2: Maximum QFI achieved via Floquet-engineered TAT-and-turn dynamics. (a) The time-evolution of the maximum QFI $F_{Q}^{\mathrm{max}}$ for different $\delta$ with the particle number $N=100$, the driving frequency $\omega=2\pi\times 10N\chi$ and the angle $\alpha=0$. (b) The time-evolution of the maximum QFI for the Floquet-engineered TAT-and-turn with $\delta_{\mathrm{opt}}$. The results of OAT, XYZ [105] and ideal TAT-and-turn are presented for comparison. (c) The enlarged area of (b) for $0\leq\chi t\leq 0.05\pi$. (d) The probability distribution $P_{m_{y}}=|\langle m_{y}|\psi\rangle|^{2}$ ($|m_{y}\rangle$ denoting the eigenstates of $\hat{J}_{y}$) and the Husimi distribution (inset) for the optimal state with largest QFI. (e) The critical detuning $\delta_{\mathrm{opt}}$ and (f) the critical evolution time $\chi t_{\mathrm{opt}}$ corresponding to (g) the optimal QFI $F_{Q}^{\mathrm{opt}}$ for different particle numbers $N$, in which the semi-classical predictions are $\delta_{\mathrm{opt}}^{\mathrm{SC}}/(N\chi)\approx 0.3135$, $t_{\mathrm{opt}}^{\mathrm{SC}}={3(1.9+0.55\ln{N})}/{N}$ and $(F_{Q}^{\mathrm{opt}})^{\mathrm{SC}}/N\approx 0.97N$, respectively. The fitted line for the optimal $F_{Q}^{\mathrm{max}}$ is $F_{Q}^{\mathrm{opt}}/N\simeq 0.77N$. (h) The optimal QFI $F_{Q}^{\mathrm{opt}}$ versus the decoherence strength $\Gamma$ in generating GHZ or GHZ-like state by OAT, XYZ, and TAT-and-turn dynamics, respectively. The black dashed (dash-dotted) line indicates the SQL (HL)
@@
Using $e^{i\gamma}=\cos{\gamma}+i\sin{\gamma}$ and $e^{iz\sin\theta}=\mathcal{{J}}_{0}(|z|)+\sum_{n=1}^{\infty}[\mathcal{{J}}_{n}(z)e^{in\theta}+(-1)^{n}\mathcal{{J}}_{n}(z)e^{-in\theta}]$ with $\mathcal{{J}}_{n}(z)$ being the Bessel function of the first kind, through performing the Fourier decomposition, the Hamiltonian (2) becomes
@@
where $\hat{H}_{0}^{I}=\frac{\chi}{2}[(1+L_{0})\hat{J}_{z}^{2}+(1-L_{0})\hat{J}_{\beta}^{2}]+K_{0}\delta\hat{J}_{z}$ and $\hat{H}_{n}^{I}$ represents $n$-order Hamiltonian with $L_{n}=\mathcal{{J}}_{n}\left(2\Omega_{0}/\omega\right)$ and $K_{n}=\mathcal{{J}}_{n}\left(\Omega_{0}/\omega\right)$ [73]. Finally, using the Floquet-Magnus expansion [94, 95] up to the order of $\omega^{-1}$, the effective Hamiltonian for Eq. (13) can be written as $\hat{H}_{\mathrm{F}}^{\rm eff}=\hat{H}_{0}^{I}+\sum_{n=1}^{\infty}\frac{1}{n\omega}([\hat{H}_{n}^{I},\hat{H}_{-n}^{I}]-[\hat{H}_{n}^{I},\hat{H}_{0}^{I}]+[\hat{H}_{-n}^{I},\hat{H}_{0}^{I}])+\mathcal{O}(\omega^{-1})$. When $\omega$ is sufficiently large ($\omega\gg N\chi$), it reduces to $\hat{H}_{\mathrm{F}}^{\rm eff}\simeq\hat{H}_{0}^{I}$ [73]. For a conserved collective spin, $\hat{J}^{2}=\hat{J}_{\alpha}^{2}+\hat{J}_{\beta}^{2}+\hat{J}_{z}^{2}=\frac{N}{2}(\frac{N}{2}+1)$ is a constant, the effective time-independent Floquet Hamiltonian is equivalent to
@@
with the effective interaction strength $\chi_{\mathrm{eff}}=\chi/3$ and the effective detuning $\delta_{\mathrm{eff}}=K_{0}\delta\simeq 0.4404\delta$. Since the TAT-and-turn Hamiltonian (5) is derived from the Hamiltonian (11) with the Floquet theorem, we refer the original time-evolution to the Floquet-engineered TAT-and-turn (labeled as FE-TATNT) dynamics. To evaluate the FE-TATNT dynamics, we also give the dynamics of the ideal TAT-and-turn Hamiltonian (5) for comparison. Our FE-TATNT dynamics contains more general features and it can be reduced to typical TAT [100, 101], OAT-and-turn, or OAT dynamics by varying system parameters [73].
Fast generation of GHZ-like states using Floquet-engineered TAT-and-turn.— The desired entangled states are generated from from an initial state $\hat{\rho}(0)=|\psi_{0}\rangle\langle\psi_{0}|$, which is easy to prepare, according to the time-evolution $\dot{\hat{\rho}}=-i[\hat{H},\hat{\rho}]$. To characterize their full metrological potential, we employ the maximum QFI $F_{Q}^{\max}$, which can be derived by computing the largest eigenvalue of the QFI matrix  [76, 77],
$\mathcal{F}\vec{n}_{\mathrm{max}}=F_{Q}^{\max}\vec{n}_{\mathrm{max}}.$ (6)
Here, the elements of the QFI matrix are expressed as
$\mathcal{F}_{\mu\nu}=\sum_{i,j=1;\kappa_{i}+\kappa_{j}\neq 0}^{\dim[\hat{\rho}]}\frac{2\mathcal{R}\left(\langle\kappa_{i}|[\hat{J}_{\mu},\hat{\rho}]|\kappa_{j}\rangle\langle\kappa_{j}|[\hat{\rho},\hat{J}_{\nu}]|\kappa_{i}\rangle\right)}{\kappa_{i}+\kappa_{j}},$ (7)
with $\mu,\nu\in\{x,y,z\}$, the spectral decomposition $\hat{\rho}=$ $\sum_{i}\kappa_{i}|\kappa_{i}\rangle\langle\kappa_{i}|$, and $\mathcal{R}$ represents the real part. The optimal signal-encoding-direction vector is $\vec{n}_{\mathrm{max}}=(n_{x}^{\mathrm{max}},n_{y}^{\mathrm{max}},n_{z}^{\mathrm{max}})^{\mathrm{T}}$, and the responding optimal generator is $\hat{J}_{\vec{n}_{\mathrm{max}}}=n_{x}^{\mathrm{max}}\hat{J}_{x}+n_{y}^{\mathrm{max}}\hat{J}_{y}+n_{z}^{\mathrm{max}}\hat{J}_{z}$. Below we use $F_{Q}^{\rm max}$ to analyze the full metrological potential of the states generated by the FE-TATNT dynamics (11).
@@
Without loss of generality, we choose $\alpha=0$ hereafter. It is advantageous to begin with an initial spin coherent state on one of the poles along $z$-axis ($\lvert\uparrow\rangle^{\otimes N}$ or $\lvert\downarrow\rangle^{\otimes N}$), which correspond to two saddle points $(0,0,\pm 1)$ in the classical phase-space of TAT-and-turn when $\left\lvert\delta_{\rm eff}/\chi_{\rm eff}\right\lvert<N$, see Fig. 1 (c). To identify the optimal condition for FE-TATNT, we analyze the dependence of $F_{Q}^{\rm max}$ on the detuning $\delta$ and the evolution time $\chi t$ in Fig. 2 (a). It clearly shows that the maximum QFI has a symmetrical distribution with respect to $\delta=0$. Given the particle number $N=100$, the maximum QFI reaches its optimal value $F_{Q}^{\mathrm{opt}}\simeq 0.77N^{2}$ when $\chi t_{\mathrm{opt}}\approx 0.13$ and $\delta_{\mathrm{opt}}\approx\pm 0.3135N\chi$. In Fig. 2 (b), we show the time-evolution of $F_{Q}^{\rm max}$ for OAT ($\hat{H}_{\rm OAT}=\chi\hat{J}_{z}^{2}$), XYZ ($\hat{H}_{\rm XYZ}=\tilde{\alpha}\chi(\hat{J}_{x}\hat{J}_{y}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{y}\hat{J}_{x})/3$ with $\tilde{\alpha}=0.4$ [105]), ideal TAT-and-turn ($\hat{H}_{\rm TATNT}$ with $\delta_{\mathrm{eff}}=K_{0}\delta_{\mathrm{opt}}$), and FE-TATNT ($\hat{H}_{\rm FE}$ with $\Omega_{0}/\omega\simeq 1.6262$ and $\delta=\delta_{\mathrm{opt}}$). The result of FE-TATNT is highly consistent with the ideal TAT-and-turn, thereby confirming the validity of the effective TAT-and-turn Hamiltonian. The FE-TATNT dynamics can attain a significantly high QFI in a very short time $\chi t\approx 0.05\pi$. In contrast, it requires a longer evolution time $\chi t\approx 0.5\pi$ for OAT or $\chi t\approx 0.1\pi$ for XYZ to reach the same level of QFI, see Figs. 2 (b) and (c). Furthermore, in the FE-TATNT with $\chi t_{\mathrm{opt}}$ and $\delta_{\mathrm{opt}}$, the generated state is a GHZ-like state [73] with two main parts distributed near two poles of the $y$ axis, which is similar to a spin cat state [32, 97, 34], see Fig. 2 (d). These results reveal that the GHZ-like state generated by FE-TATNT is much faster than that of using OAT or effective three-body interaction.
@@
The optimal QFI, the critical detuning and the critical evolution time can be found analytically using semi-classical treatment [102, 96]. Similarly to the spin cat state in the inset of Fig. 2 (d), whose variance is approximately stretched along the $y$-axis, one can estimate the global optimal parameter $\delta_{\rm opt}^{\rm SC}=\pm(\sqrt{2}-1)N\chi/(3K_{0})\approx\pm 0.3135N\chi$ using the upper boundary of the QFI [73]. Due to the fact that the center of the uncertainty patch is fixed, the TAT-and-turn dynamics merely alters the distribution by stretching and squeezing. Thus, the relevant timescale can be estimated by the time interval between the boundary points of the uncertainty patch along the separatrix [96], which can be given as [73]
@@
Based on the above analysis, the best achievable QFI exhibits the Heisenberg scaling, which matches the ultimate precision bound of the spin cat states [73]. As shown in Figs. 2 (e)-(g), the above results are in accordance with the numerical results of FE-TATNT.
According to Eq. (8), the TAT-and-turn dynamics exhibits a significantly shorter timescale for large particle numbers, which is advantageous in mitigating the decoherence effect. Here we consider the decoherence effect of spontaneous emission. The time-evolution can be described by [73] $\dot{\hat{\rho}}=-i[\hat{H},\hat{\rho}]+\left(\hat{L}\hat{\rho}\hat{L}^{\dagger}-\frac{1}{2}\hat{L}^{\dagger}\hat{L}\hat{\rho}-\frac{1}{2}\hat{\rho}\hat{L}^{\dagger}\hat{L}\right)$ where the Lindblad operator is $\hat{L}=\sqrt{\Gamma/2}\hat{J}_{-}$ with $\hat{J}_{-}=\hat{J}_{x}-i\hat{J}_{y}$ and the spontaneous emission rate $\Gamma$ [106, 39, 105]. As shown in Fig. 2 (h), we respectively employ OAT, XYZ and TATNT dynamics to generate GHZ or GHZ-like state under decoherence, showing that the TATNT protocol is more robust against decoherence than the OAT and XYZ protocols, and demonstrating remarkable metrological usefulness in quantum metrology with noises. Similar results for correlated dephasing are also shown in the Supplementary Material [73].
@@
Figure 3: (a) Measurement precision yielded from Floquet-engineered anti-TAT-and-turn. The metrological gain is defined as $G=20\log_{10}\left[{(\Delta\phi)_{\mathrm{SQL}}}/{\Delta\phi}\right]$ with $(\Delta\phi)_{\mathrm{SQL}}=1/{\sqrt{N}}$ denoting the measurement precision of SQL. (b) Robustness against detection noise. The green dashed line is the result of population measurement $\langle\hat{J}_{z}\rangle_{\sigma}$ after Floquet-engineered anti-TAT-and-turn with $\chi t=0.12$. The blue dash-dotted and yellow dotted lines are results of parity measurement ^$\langle\hat{\prod}\rangle_{\sigma}$ for GHZ-like and GHZ states generated by XYZ and OAT dynamics, respectively. Here, the black dashed (dash-dotted) line indicates the SQL (HL). The driving frequency is set to $\omega=2\pi\times 100N\chi$, and $\alpha=\pi/2$. These parameters, along with a particle number of $N=100$ and an estimated phase of $\phi=1/1000$, are selected for the simulation of FE-TATNT.